The mean free path of photons in stars is of the order of 0.2 cm. Show that the time
taken for a photon to reach the surface of a star of radius 4 Ro is of the order of one
million year.
1
Expert's answer
2017-02-23T13:14:05-0500
The time for beam traveling is given by the expression t = (Ro)^2/(l*c) or in our case t = 16*(Ro)^2/(l*c). Thus, t = 16*(6.96*10^8 m)^2/(2*10^-3 m * 3*10^8 m/s) = 412000 years.